OpenStudy (anonymous):

What is the correct factored form of the given equation? 4x^2 - 25 = 0 2(x - 5)² = 0?

OpenStudy (lgbasallote):

not exactly...this is difference of two squares

mathslover (mathslover):

\[\large{4x^2-25=0}\] \[\large{(2x)^2-(5)^2=0}\] right?

OpenStudy (lgbasallote):

\[\huge a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

(2x - 5)(2x + 5) = 0

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

like this?

mathslover (mathslover):

very good @browniesmiths98

OpenStudy (lgbasallote):

yes exactly

mathslover (mathslover):

good work ..

OpenStudy (anonymous):

thankyou:)

mathslover (mathslover):

your welcome :) and : \[\huge{\mathbb{Welcome}\textbf{To}\mathbb{Openstudy}}\]

OpenStudy (anonymous):

:) yaaaa